# Tiny models from tiny data: Textual and null-text inversion for few-shot distillation

### v1

@@ Abstract
Few-shot image classification involves classifying images using very few training examples. Recent vision foundation models show excellent few-shot transfer abilities, but are large and slow at inference. Using knowledge distillation, the capabilities of high-performing but slow models can be transferred to tiny, efficient models. However, common distillation methods require a large set of unlabeled data, which is not available in the few-shot setting. To overcome this lack of data, there has been a recent interest in using synthetic data. We expand on this work by presenting a novel diffusion model inversion technique (TINT) combining the diversity of textual inversion with the specificity of null-text inversion. Using this method in a few-shot distillation pipeline leads to state-of-the-art accuracy among small student models on popular benchmarks, while being significantly faster than prior work. This allows us to push even tiny models to high accuracy using only a tiny application-specific dataset, albeit relying on extra data for pre-training.
Popular few-shot benchmarks involve evaluation over a large number of episodes, which is computationally cumbersome for methods involving synthetic data generation. Therefore, we also present a theoretical analysis on how the variance of the accuracy estimator depends on the number of episodes and query examples, and use these results to lower the computational effort required for method evaluation. In addition, to further motivate the use of generative models in few-shot distillation, we demonstrate that our method performs better compared to training on real data mined from the dataset used to train the diffusion model.
Source code will be made available at https://github.com/pixwse/tiny2.
@@ 1 Introduction
In recent years, deep learning has enabled unprecedented performance in most computer vision tasks. To reach state-of-the-art, increasingly large foundation models [1, 2, 3, 4, 5] are often used, trained on increasingly huge datasets [6, 7]. In contrast, many computer vision applications have the opposite requirements. Representative training data can be scarce due to privacy concerns (e.g. medical data) or other practical issues around data collection, and lightning-fast inference is desired in both embedded systems and to get good economy in cloud-hosted solutions. However, many applications have no restrictions on using large pretrained models in the training pipelines, as long as the final model is small and efficient. Our problem statement can thus be formulated as: How can we maximize the accuracy of tiny classifiers if only a tiny amount of application-specific labeled data is available (few-shot setting), but there are no restrictions on using common pre-trained models in the training pipeline?
@@
Large foundation models often present good few-shot transfer abilities when equipped with a small classifier head [8, 9, 10]. While large foundation models may be required to handle any few-shot task well, a specific task (containing just a few classes) is likely solvable using a significantly smaller model. Using knowledge distillation [11, 12], a small, efficient classifier could be trained to mimic a large general few-shot model on a specific task. However, while typical distillation methods require a large set of representative unlabeled data [11, 13, 14], we only assume access to a few application-specific examples. This leads us to our overall approach: Adapt a generative model to produce a large set of images representative of the novel classes. Then use this data to distill knowledge from a well-performing (but large and slow) few-shot classifier to a small model, specialized on a single task. We base the generator on diffusion models, due to their impressive image generation capabilities [15, 16, 17, 18]. The core challenge is how to best specialize such models to the novel few-shot classes. To this end, we propose a novel combination of textual inversion and null-text inversion that is illustrated in Figure 1, with some examples of generated training images shown in Figure 2.
@@
Common few-shot benchmarks such as miniImageNet are often strictly interpreted in the sense that only training data from within the dataset itself may be used. In their recent work [8], Hu et al. argue that maintaining this strict interpretation may lead to an undesirable divergence between the few-shot learning and semi-supervised learning [19] communities. We adhere to this view, as our goal is to present a pipeline that is useful for real-world applications without any restrictions on additional data sources.
Our contributions can be summarized as:
  • A novel diffusion inversion method (TINT) combining the diversity of textual inversion with the specificity of null-text inversion.
  • Reaching state-of-the-art accuracy on miniImageNet, most notably significantly exceeding the best previously reported accuracy for the tiny Conv4 model, while reducing the training time compared to competing methods with an order of magnitude.
  • A theoretical analysis of the accuracy estimator variance in episode-based benchmarks, enabling faster evaluation methods without reduced variance.
  • A demonstration that using a generative model in few-shot distillation outperforms using the original training data (LAION [6]) of the generative model directly.
@@ 2.1 Few-shot classification
In few-shot classification [20, 21], the task is to classify a query $\mathbf{x}^{\textrm{(q)}}$ into one of $N$ classes identified by integer labels $y\in\{1,...,N\}$, where each class is defined using a small set of labeled support examples. To simplify the notation, we focus on the case with $K$ support examples per class (N-way, K-shot), and use $\mathbf{x}^{\textrm{(s)}}_{n,k}$ to denote the $k$’th support example of class $n$. All theory transfers naturally to the case where $K$ is different for each class.
@@ 2.2 Diffusion Models
Denoising diffusion probabilistic models [22] model the distribution of a random variable $\mathbf{x}_{0}$ by transforming it into a tractable distribution (noise) over timesteps $t\in\{0,...,T\}$. Our generator is built upon the popular Stable Diffusion (SD) [15] model due to its public availability and use in prior work on diffusion inversion [23, 24]. SD defines the diffusion process on a latent variable $\mathbf{z}_{t}$ of lower dimensionality than $\mathbf{x}_{t}$. The model is trained to minimize the loss
@@
where the noise estimation $\epsilon(\mathbf{z}_{t};t,\mathbf{v})$ is implemented using a U-net [25], and $\mathbf{v}$ is a CLIP [26] embedding of a text prompt fed as an additional input to this U-net. A VQ-VAE model [27] maps between $\mathbf{x}_{0}$ and $\mathbf{z}_{0}$.
Using Denoising Diffusion Implicit Models (DDIM) [28], the forward and backward processes can be made deterministic. We write this as $\mathbf{z}_{t+1}=f_{t}(\mathbf{z}_{t})$, $\mathbf{z}_{t-1}=g_{t}(\mathbf{z}_{t})$. We are mostly interested in the mapping between $\mathbf{z}_{0}$ and $\mathbf{z}_{T}$, writing it as $\mathbf{z}_{T}=F(\mathbf{z}_{0})$ with $F=f_{T-1}\circ...\circ f_{0}$ and $\mathbf{z}_{0}=G(\mathbf{z}_{T})$ with $G=g_{1}\circ...\circ g_{T}$. The strength of the conditioning input can be controlled by classifier-free guidance (CFG) [29]. For each timestep $t$, the noise is predicted using both a text embedding $\mathbf{v}$ and an unconditional embedding $\mathbf{u}$. The noise prediction is then shifted in the direction produced by $\mathbf{v}$ according to
@@
When using classifier-free guidance, the DDIM inversion no longer reproduces the original input unless additional techniques are used. Null-text inversion (NTI) [24] is one such technique. In NTI, latents $\mathbf{z}_{T}=F(\mathbf{z}_{0})$ are first computed using DDIM. The original objective (Eq. 1) is then minimized with the unconditional embeddings $\mathbf{u}_{1:T}$ as variables (introducing a separate embedding $\mathbf{u}_{t}$ for each timestep $t$). The output of the inversion is the final latent $\mathbf{z}_{T}$ and adjusted embeddings $\mathbf{u}_{1:T}$.
Textual inversion (TI) [23] is a technique for making diffusion models output images that look similar (but not identical) to a small set of input images. The idea is to adjust the text embeddings of one or more words in a conditioning text prompt, essentially creating new words representing the input images. The optimization objective is still the original Eq. 1, but the variables to optimize are now selected embeddings of individual words (parts of $\mathbf{v}$).
@@ 2.3 Knowledge distillation
Knowledge distillation [30, 11, 12] refers to training a student network to mimic a teacher network for the purpose of model compression or performance improvement. We base our pipeline on the LabelHalluc method [31], where distillation is used in few-shot classification, using base classes as distillation training data, pseudo-labelled by a teacher sharing the same architecture as the student.
@@ 3.1 TINT
The method is illustrated in Figure 1 and works as follows: First, all support examples $\mathbf{x}^{\textrm{(s)}}_{n,k}$ of class $n$ are fed jointly to a TI procedure to find a text embedding $\mathbf{v}_{n}$ capable of generating similar examples. We use the generic prompt a photo of an #object, where #object represents the embedding to optimize. The embedding is initialized from the word object, i.e. we don’t require a known text description of each support class. We then run NTI on each support example separately, conditioned by the embeddings obtained from TI. This produces latent vectors $\mathbf{z}_{n,k}$ at time $t=T$ and adjusted unconditional embeddings $\mathbf{u}_{n,1:T,k}$. Note that we have now optimized all embeddings present in the CFG procedure (Eq. 2) - both the conditional and unconditional ones.
@@
Given the optimized embeddings, we want to make a random perturbation of a latent $\mathbf{z}_{n,k}$ and feed it back through the generative model $G$, using the adjusted $\mathbf{u}_{n,1:T,k}$ in the CFG. A naive way to do the random perturbation would be to blend $\mathbf{z}_{n,k}$ with noise, i.e. letting
$\mathbf{z}=(1-\alpha)\mathbf{z}_{n,k}+\alpha\mathbf{n},\quad\mathbf{n}\sim\mathcal{N}(0,\mathbf{I}),\quad\alpha\in[0,1].$ (3)
However, as noted in [32], a simple linear interpolation between two latents will often result in an output with lower norm than the inputs and lead to an overly bland, low-contrast image. To avoid this effect, we could simply rescale the latent norm by letting $\mathbf{\tilde{z}}=\mathbf{z}\|\mathbf{z}\|^{-1}$, but that would not reproduce $\mathbf{z}_{n,k}$ as $\alpha\rightarrow 0$. One option could be to use the norm-guided interpolation suggested by [32], but we opt for a simpler approach. We compute a desired latent norm by interpolating the norms of the inputs to the blending operation from Eq. 3 and rescaling $\mathbf{z}$ to this norm value, by
$\mathbf{\tilde{z}}=\mathbf{z}\|\mathbf{z}\|^{-1}\big{(}(1-\alpha)\|\mathbf{z}_{n,k}\|+\alpha\|\mathbf{n}\|\big{)}.$ (4)
@@
Note that there is a significant resolution mismatch between available pre-trained Stable Diffusion models and popular few-shot classification datasets. We found that a naive upsampling before the textual inversion made the inversion prone to reproduce sampling artifacts, sometimes prioritizing this over semantically meaningful details. Using a pre-trained super-resolution (SR) model [33] for upscaling largely eliminated this problem. We also found it beneficial to apply the TI optimization target (Eq. 1) only for a restricted range $t\in[250,1000]$, as including smaller $t$ made the textual inversion more prone to focus on remaining sampling artifacts. See Appendix B.2 for more details.
@@ 3.2 Distillation
The distillation pipeline is illustrated in Figure 4. It largely follows prior work [31, 34, 32], but uses a strong off-the-shelf teacher and our novel image generation method. As noted in [31], images from the base classes (meta-training split) of a few-shot dataset can be pseudo-labeled by a teacher and used for distillation. We include this idea, leaving us with 3 types of distillation training data: (1) training images from base classes, (2) TINT-generated synthetic images of novel classes, and (3) support images from novel classes. The base and synthetic images are pseudo-labeled using a strong off-the-shelf teacher. See Appendix B.4 for more details.
@@ 3.3 Efficient multi-episode evaluation
Few-shot classifiers are often evaluated over randomly drawn episodes, where each episode contains a random set of novel classes with random support and query examples. For example, results on miniImageNet [20] are typically evaluated over 600 episodes, each containing 5 classes with $K$ (1 or 5) support examples and $Q=15$ query examples per class. In our pipeline, specialization involves generating thousands of synthetic images, and repeating this process 600 times or more is computationally cumbersome. However, evaluating over a larger query set would use negligible compute in comparison. Instead of the most common choice of 15 query examples per class, we could easily use the remaining 595 queries per class for miniImageNet.
A question then arises: can we trade episodes for query examples, i.e. run the evaluation over fewer episodes but with more query examples per episode, keeping a similar statistical significance in the result? As a guide for what happens when varying the number of episodes and queries, we present the following theorem:
@@
Let $P,Q\in\mathbb{Z}^{+}$ be the number of episodes and query examples per episode respectively. Let $a_{p}\in[0,1]$ be the true (unknown) accuracy of an evaluated method on episode $p\in\{1,...,P\}$ and let $a=\mathbb{E}[a_{p}]$ and $\sigma_{a}^{2}=\mathbb{V}[a_{p}]$ be the true (unknown) mean and variance of $a_{p}$ over i.i.d. episodes $p$.
@@
$\mathbb{V}[\tilde{a}]$ $=\frac{1}{P}\left(\frac{1}{Q}a(1-a)+\left(1-\frac{1}{Q}\right)\sigma_{a}^{2}\right).$ (6)
Eq. 5 means that the estimated accuracy $\tilde{a}$ is an unbiased estimate of the true unknown $a$ regardless of $P$ and $Q$. Eq. 6 shows how $P$ and $Q$ affect the estimator variance, and can be used e.g. to determine suitable values for $P$ and $Q$ to reach a certain desired estimator variance. To arrive at Eq. 6, we model the evaluation over each episode as a set of Bernoulli trials and derive an expression of $\mathbb{V}[\tilde{a}]$ using the law of total variance. Motivated by this relation, we decided to reduce $P$ to 120 while increasing $P$ to $595N$ in our evaluation. The full proof and a numerical example motivating this choice can be found in Appendix A.3.

### v2

@@ Abstract
Few-shot learning deals with problems such as image classification using very few training examples. Recent vision foundation models show excellent few-shot transfer abilities, but are large and slow at inference. Using knowledge distillation, the capabilities of high-performing but slow models can be transferred to tiny, efficient models. However, common distillation methods require a large set of unlabeled data, which is not available in the few-shot setting. To overcome this lack of data, there has been a recent interest in using synthetic data. We expand on this line of research by presenting a novel diffusion model inversion technique (TINT) combining the diversity of textual inversion with the specificity of null-text inversion. Using this method in a few-shot distillation pipeline leads to state-of-the-art accuracy among small student models on popular benchmarks, while being significantly faster than prior work. Popular few-shot benchmarks involve evaluation over a large number of episodes, which is computationally cumbersome for methods involving synthetic data generation. We also present a theoretical analysis on how the accuracy estimator variance depends on the number of episodes and query examples, and use these results to lower the computational effort required for method evaluation. Finally, to further motivate the use of generative models in few-shot distillation, we demonstrate that our method outperforms training on real data mined from the dataset used in the original diffusion model training. Source code is available at https://github.com/pixwse/tiny2.
@@ 1 Introduction
In recent years, deep learning has enabled unprecedented performance in most computer vision tasks. To reach state-of-the-art, increasingly large foundation models (Dosovitskiy et al., 2020; Touvron et al., 2021; Fang et al., 2023; Kirillov et al., 2023; Awais et al., 2023) are often used, trained on increasingly huge datasets (Schuhmann et al., 2022; Zhai et al., 2022). In contrast, many computer vision applications have the opposite requirements. Representative training data can be scarce due to privacy concerns (e.g. medical data) or other practical issues around data collection (e.g. industrial fault detection requiring destructive processes for obtaining training data). Lightning-fast inference is often desired in both embedded systems and to get good economy in cloud-hosted solutions.
Few-shot learning (Song et al., 2023) deals with learning in very label-constrained situations, typically down to 1-5 labelled examples per class. Classical few-shot methods study architectures that can quickly adapt to new tasks by feeding a new set of labelled support examples into the model. Common few-shot benchmarks such as miniImageNet (Vinyals et al., 2016) have traditionally been strictly interpreted in the sense that only training data from within the dataset itself may be used, but models trained on extra data are becoming more and more common (Samuel et al., 2024a; Trabucco et al., 2024). Hu et al. (2022) argue that the strict (no pretraining) interpretation of few-shot learning may lead to an undesirable divergence between the few-shot learning and semi-supervised learning (Van Engelen & Hoos, 2020) communities. We adhere to this view.
In many practical applications, longer training times for adaptation to a specific task and the use of pre-trained models are often both acceptable, as long as the final model for a specific task has excellent performance in terms of speed and accuracy. To address this setting, we formulate our research question as: How can we maximize the accuracy of tiny efficient classifiers when only a tiny amount of application-specific labeled data is available, but when there are no restrictions on using common pre-trained models in the training pipeline, and a few GPU-hours of training for a specific task is acceptable? We emphasize that this is not an artificial scenario - rather, it is motivated from personal experience from MVP development of new industrial applications.
@@
Large foundation models often present good few-shot transfer abilities when equipped with a small classifier head (Hu et al., 2022; Caron et al., 2021; Zhang et al., 2022). While such models may be required to handle any few-shot task well, a specific task (containing just a few classes) is likely solvable using a significantly smaller model. Using knowledge distillation (Hinton et al., 2015; Gou et al., 2021), a small, efficient classifier can be trained to mimic a large general few-shot model on a specific task. However, while typical distillation methods require a large set of representative unlabeled data (Hinton et al., 2015; Tian et al., 2019; Ahn et al., 2019), we only assume access to a few application-specific examples. This leads us to our overall approach: Adapt a generative model to produce a large set of images representative of the novel classes. Then use this data to distill knowledge from a well-performing (but large and slow) few-shot classifier to a small model, specialized on a single task. We base the generator on diffusion models, due to their impressive image generation capabilities (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; Balaji et al., 2022). A core challenge then becomes how to best specialize such models to produce data similar to the few-shot support examples. To this end, we propose a novel combination of textual inversion and null-text inversion, illustrated in Figure 1, and with some examples of generated training images shown in Figure 2.
@@
Evaluation on traditional few-shot datasets (e.g. miniImageNet) is often done over a large number of randomly drawn episodes. This poses a practical problem for methods where specialization for each episode requires significant computational effort. In Section 3.3, we analyze the impact of the number of episodes and number of query examples per episode, and arrive at an evaluation setup reaching similar statistical significance with less compute. This makes it more practical to evaluate methods with heavy specialization procedures on traditional few-shot datasets.
Our main contributions can be summarized as:
  • A novel diffusion inversion method (TINT), combining the diversity of textual inversion with the specificity of null-text inversion and outperforming baseline generation methods in the few-shot distillation context.
  • Demonstrating that it is possible to significantly push the achievable accuracy of even tiny models such as the 113k-parameter Conv4 by using them in a generative distillation framework.
  • A theoretical analysis of the accuracy estimator variance in episode-based few-shot benchmarks, arriving at an evaluation procedure that is more practical for specialization-heavy methods.
  • Demonstrating that using a generative model in few-shot distillation outperforms a direct use of the LAION dataset (Schuhmann et al., 2022) which was used to train the generative model in the first place.
@@ 2.1 Few-shot classification
In few-shot classification (Vinyals et al., 2016; Ravi & Larochelle, 2016), the task is to classify a query $\mathbf{x}^{\textrm{(q)}}$ into one of $N$ classes identified by integer labels $y\in\{1,...,N\}$, where each class is defined using a small set of labeled support examples. To simplify the notation, we focus on the case with $K$ support examples per class (N-way, K-shot), and use $\mathbf{x}^{\textrm{(s)}}_{n,k}$ to denote the $k$’th support example of class $n$. All theory transfers naturally to the case where $K$ is different for each class.
@@ 2.2 Diffusion Models
Denoising diffusion probabilistic models (Ho et al., 2020) model the distribution of a random variable $\mathbf{x}_{0}$ by transforming it into a tractable distribution (noise) over timesteps $t\in\{0,...,T\}$. Our generator is built upon the popular Stable Diffusion (SD) (Rombach et al., 2022) model due to its public availability and use in prior work on diffusion inversion (Gal et al., 2022; Mokady et al., 2023). SD defines the diffusion process on a latent variable $\mathbf{z}_{t}$ of lower dimensionality than $\mathbf{x}_{t}$. The model is trained to minimize the loss
@@
where the noise estimation $\epsilon(\mathbf{z}_{t};t,\mathbf{v})$ is implemented using a U-net (Ronneberger et al., 2015), and $\mathbf{v}$ is a CLIP embedding (Radford et al., 2021) of a text prompt fed as an additional input to this U-net. A VQ-VAE model (Razavi et al., 2019) maps between $\mathbf{x}_{0}$ and $\mathbf{z}_{0}$.
Using Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020), the forward and backward processes can be made deterministic. We write this as $\mathbf{z}_{t+1}=f_{t}(\mathbf{z}_{t})$, $\mathbf{z}_{t-1}=g_{t}(\mathbf{z}_{t})$. We are mostly interested in the mapping between $\mathbf{z}_{0}$ and $\mathbf{z}_{T}$, writing it as $\mathbf{z}_{T}=F(\mathbf{z}_{0})$ with $F=f_{T-1}\circ...\circ f_{0}$ and $\mathbf{z}_{0}=G(\mathbf{z}_{T})$ with $G=g_{1}\circ...\circ g_{T}$. The strength of the conditioning can be controlled by classifier-free guidance (CFG) (Ho & Salimans, 2022). For each timestep $t$, the noise is predicted using both a text embedding $\mathbf{v}$ and an unconditional embedding $\mathbf{u}$, and the noise prediction is shifted in the text-conditional direction according to
@@
When using classifier-free guidance, the DDIM inversion no longer reproduces the original input unless additional techniques are used. Null-text inversion (NTI) (Mokady et al., 2023) is one such technique. In NTI, latents $\mathbf{z}_{T}=F(\mathbf{z}_{0})$ are first computed using DDIM. The original objective (Eq. 1) is then minimized with the unconditional embeddings $\mathbf{u}_{1:T}$ as variables (introducing a separate embedding $\mathbf{u}_{t}$ for each timestep $t$). The output of the inversion is the final latent $\mathbf{z}_{T}$ and adjusted embeddings $\mathbf{u}_{1:T}$.
Textual inversion (TI) (Gal et al., 2022) is a technique for making diffusion models output images that look similar (but not identical) to a small set of input images. The idea is to adjust the text embeddings of one or more words in a conditioning text prompt, essentially creating new words representing the input images. The optimization objective is still the original Eq. 1, but the variables to optimize are now selected embeddings of individual words (parts of $\mathbf{v}$).
@@ 2.3 Knowledge distillation
Knowledge distillation (Buciluǎ et al., 2006; Hinton et al., 2015; Gou et al., 2021) refers to training a student network to mimic a teacher network for the purpose of model compression or performance improvement. We base our pipeline on the LabelHalluc method Jian & Torresani (2022), where distillation is used in few-shot classification, using base classes as distillation training data, pseudo-labeled by a teacher sharing the same architecture as the student.
@@ 3.1 TINT
The method is illustrated in Figure 1 and works as follows: First, all support examples $\mathbf{x}^{\textrm{(s)}}_{n,k}$ of class $n$ are fed jointly to a TI procedure to find a text embedding $\mathbf{v}_{n}$ capable of generating similar examples. We use the generic prompt a photo of an #object, where #object represents the embedding to optimize. The embedding is initialized from the word object, i.e. we don’t require a known text description of each support class. We then run NTI on each support example separately, conditioned by the embeddings obtained from TI. This produces latent vectors $\mathbf{z}_{T,n,k}$ at time $t=T$ and adjusted unconditional embeddings $\mathbf{u}_{1:T,n,k}$. Note that we have now optimized all embeddings present in the CFG procedure (Eq. 2) - both the conditional and unconditional ones.
@@
Given the optimized embeddings, we want to make a random perturbation of a latent $\mathbf{z}_{T,n,k}$ and feed it back through the generative model $G$, using the adjusted $\mathbf{u}_{1:T,n,k}$ in the CFG. A naive way to do the random perturbation would be to blend $\mathbf{z}_{T,n,k}$ with noise, i.e. letting
$\mathbf{z}^{\prime}=(1-\alpha)\mathbf{z}_{T,n,k}+\alpha\mathbf{n},\quad\mathbf{n}\sim\mathcal{N}(0,\mathbf{I}),\quad\alpha\in[0,1].$ (3)
However, as noted by Samuel et al. (2024a), a simple linear interpolation between two latents will often result in an output with lower norm than the inputs and lead to an overly bland, low-contrast image. To avoid this effect, we could simply rescale the latent norm by letting $\mathbf{\tilde{z}}=\mathbf{z}^{\prime}\|\mathbf{z}^{\prime}\|^{-1}$, but that would not reproduce $\mathbf{z}_{T,n,k}$ as $\alpha\rightarrow 0$. One option could be to use the norm-guided interpolation suggested by Samuel et al. (2024a), but we opt for a simpler approach. We compute a desired latent norm by interpolating the norms of the inputs to the blending operation from Eq. 3 and rescaling $\mathbf{z}^{\prime}$ to this norm value, by
$\mathbf{\tilde{z}}=\mathbf{z}^{\prime}\|\mathbf{z}^{\prime}\|^{-1}\big{(}(1-\alpha)\|\mathbf{z}_{T,n,k}\|+\alpha\|\mathbf{n}\|\big{)}.$ (4)
@@
Note that there is a significant resolution mismatch between available pre-trained Stable Diffusion models and popular few-shot classification datasets. We found that a naive upsampling before the textual inversion made the inversion prone to reproduce sampling artifacts, sometimes prioritizing this over semantically meaningful details. Using a pre-trained super-resolution (SR) model (Chen et al., 2023) for upscaling largely eliminated this problem. We also found it beneficial to apply the TI optimization target (Eq. 1) only for a restricted range $t\in[250,1000]$, as including smaller $t$ made the textual inversion more prone to focus on remaining sampling artifacts. See Appendix B.2 for more details.
@@ 3.2 Distillation
The distillation pipeline is illustrated in Figure 4. It largely follows prior work (Jian & Torresani, 2022; Roy et al., 2022; Samuel et al., 2024a), but uses a strong off-the-shelf teacher and our novel image generation method. As noted in Jian & Torresani (2022), images from the base classes (meta-training split) of a few-shot dataset can be pseudo-labeled by a teacher and used for distillation. We include this idea, leaving us with 3 types of distillation training data: (1) training images from base classes, (2) TINT-generated synthetic images of novel classes, and (3) support images from novel classes. The base and synthetic images are pseudo-labeled using a strong off-the-shelf teacher. See Appendix B.4 for more details.
@@ 3.3 Efficient multi-episode evaluation
Few-shot classifiers are often evaluated over randomly drawn episodes, where each episode contains a random set of novel classes with random support and query examples. For example, results on miniImageNet (Vinyals et al., 2016) are typically evaluated over 600 episodes, each containing 5 classes with $K$ (1 or 5) support examples and $Q=15$ query examples per class. In our pipeline, specialization involves generating thousands of synthetic images, and repeating this process 600 times or more is computationally cumbersome. However, evaluating over a larger query set would use negligible compute in comparison. Instead of the traditional choice of 15 query examples per class, we could easily use the remaining 595 queries per class for miniImageNet. A question then arises: can we trade episodes for query examples, i.e. run the evaluation over fewer episodes but with more query examples per episode? How would this affect the statistical significance of the estimate? As a guide for what happens when varying the number of episodes and queries, we present the following theorem:
@@
Let $P,Q\in\mathbb{Z}^{+}$ be the number of episodes and query examples per episode respectively. Let $a_{p}\in[0,1]$ be the true (unknown) accuracy of an evaluated method on episode $p\in\{1,...,P\}$ and let $a=\mathbb{E}[a_{p}]$ and $\sigma_{a}^{2}=\mathrm{Var}[a_{p}]$ be the true (unknown) mean and variance of $a_{p}$ over i.i.d. episodes $p$.
@@
$\mathrm{Var}[\tilde{a}]$ $=\frac{1}{P}\left(\frac{1}{Q}a(1-a)+\left(1-\frac{1}{Q}\right)\sigma_{a}^{2}\right).$ (6)
Eq. 5 means that the estimated accuracy $\tilde{a}$ is an unbiased estimate of the true unknown $a$ regardless of $P$ and $Q$. Eq. 6 shows how $P$ and $Q$ affect the estimator variance, and can be used e.g. to determine suitable values for $P$ and $Q$ to reach a certain desired estimator variance. To arrive at Eq. 6, we model the evaluation over each episode as performing $Q$ Bernoulli trials with success rate $a_{p}$, which makes $\tilde{a}_{p}$ the mean of $Q$ Bernoulli-distributed variables. We can then derive an expression of $\mathrm{Var}[\tilde{a}]$ using the law of total variance. The full proof is presented in the appendix, Section A.3.
Theorem 1 allows us to study the expected impact of varying $Q$ and $P$. Note that for moderately large $Q$, Eq. 15 can be approximated as
$\mathrm{Var}[\tilde{a}]=\frac{1}{P}\left(\frac{1}{Q}a(1-a)+\sigma_{a}^{2}\right).$ (7)
As $Q$ increases, $\mathrm{Var}[\tilde{a}]$ approaches $P^{-1}\sigma_{a}^{2}$, representing the case where the accuracy of each episode is estimated perfectly, and only the inter-episode variance remains.
Eq. 15 (or Eq. 7) can be used for experiment planning, assessing the expected variance of an evaluation using a given $P$ and $Q$, or giving guidance around the number of evaluation episodes required for demonstrating a certain improvement with a desired statistical significance. As a concrete example, for the LabelHalluc method (Jian & Torresani, 2022), $a\approx 0.87,\sigma\approx 0.05$, giving an estimator variance of 7e-6 over 600 episodes using 25 query examples (5 per class). In preliminary experiments, the accuracy of our method was $a\approx 0.93$ with $\sigma\approx 0.028$. Increasing $P$ to include all remaining examples ($595N$), asserting that we want a similar estimator variance as the original LabelHalluc method and solving for $P$ gives us $P=118$. We therefore settled for 120 evaluation episodes as a reasonable trade-off between computational requirements and confidence. As can be seen in Table 5, our 95% confidence interval for ResNet12 on miniImageNet is indeed similar to that of the LabelHalluc method (0.50 vs 0.48), showing that the practical results are consistent with the theory.